IF 3 times a number is added to 4 the result is 22. How much is 1 less than twice the number?
Answer by stanbon(75887)

IF 3 times a number is added to 4 the result is 22.
3x+4 = 22
3x = 18
x = 6
----
How much is 1 less than twice the number?
2x-1 = 2*6-1 = 11
======================
Cheers,
Stan H.